# FedP2EFT: Federated Learning to Personalize Parameter Efficient Fine-Tuning for Multilingual LLMs

### v1

@@ Abstract
Federated learning (FL) has enabled the training of multilingual large language models (LLMs) on diverse and decentralized multilingual data, especially on low-resource languages. To improve client-specific performance, personalization via the use of parameter-efficient fine-tuning (PEFT) modules such as LoRA is common. This involves a personalization strategy (PS), such as the design of the PEFT adapter structures (e.g., in which layers to add LoRAs and what ranks) and choice of hyperparameters (e.g., learning rates) for fine-tuning. Instead of manual PS configuration, we propose FedP2EFT, a federated learning-to-personalize method for multilingual LLMs in cross-device FL settings. Unlike most existing PEFT structure selection methods, which are prone to overfitting low-data regimes, FedP2EFT collaboratively learns the optimal personalized PEFT structure for each client via Bayesian sparse rank selection. Evaluations on both simulated and real-world multilingual FL benchmarks demonstrate that FedP2EFT largely outperforms existing personalized fine-tuning methods, while complementing a range of existing FL methods.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: (a) FedP2EFT’s inference stage on a single client. 1 Given the base model and the client’s train dataset, features are extracted and passed into our PS generator (PSG) to generate a PS, $\bm{\lambda}$, suited to the client’s resource budget. 2 $\bm{\lambda}$ is then used to initialize all LoRA modules before 3 the base model is personalized. The resulting personalized model is then used to evaluate on the client’s test samples. (b) We train PSG using standard FL approaches.
Federated learning (FL) makes it possible to train multilingual large language models (LLMs) across different geographical regions, protecting linguistic diversity for low-resource languages (Zhao et al., 2024) while being compliant with privacy regulations (Lim et al., 2020), e.g., General Data Protection Regulation (GDPR). Despite the impressive capabilities demonstrated by these models across various languages, their performance can vary significantly depending on the specific language (Rust et al., 2021) and the data quantity (Adelani et al., 2021) on each client.
Moreover, the majority of existing FL-based multilingual LLM approaches have thus far focused on training a single global model (Iacob et al., 2025; Weller et al., 2022; Ye et al., 2024), limiting their performance on specific languages. Concretely, scaling a single model to different languages is challenged by issues such as the Curse of Multilinguality (Conneau et al., 2020), where adding more languages often leads to diminishing returns, and Negative Interference (Wang et al., 2020b), where diverse languages compete for limited model capacity. From a personalized FL perspective, learning a global model often increases the initial performance at the expense of personalized performance, e.g., fine-tuning from the global model (Jiang et al., 2019).
Naturally, personalized FL approaches can help bridge the gap and improve language personalization. However, existing techniques are either too costly to be applied to LLMs, e.g., the use of meta-learning (Chen et al., 2018) and hypernetworks (Shamsian et al., 2021), or rely on suboptimal hand-crafted personalization strategies, e.g., manual choice of personalized and language-specific layers (Iacob et al., 2025; Zhao et al., 2024; Wu et al., 2024; Qi et al., 2024), parameter-efficient fine-tuning (PEFT) adapter structures (Yang et al., 2024a, b), or clustering based on class labels and/or language commonality (Mansour et al., 2020; Sattler et al., 2020; Ye et al., 2024).
Intuitively, optimizing personalization in personalized FL often necessitates dataset- and task-specific methods. The optimal level of personalization varies significantly depending on the characteristics of the data and the specific FL scenario (Chen et al., 2022; Ye et al., 2024). For instance, an English-pretrained LLM may require stronger personalization, e.g., higher learning rates, when fine-tuning on German than on English. The optimal personalization strategy (PS) is thus contingent upon the specific task, the client, and the given base model (Lee et al., 2023).
In this paper, we address the issues above by learning to personalize PEFT for each client. To this end, we propose FedP2EFT, a method that enables clients to collaboratively learn language personalization strategies using FL. Specifically, we federatedly train a PS generator (PSG), as depicted in Fig. 1(b), which allows all participating clients to collaboratively learn the optimal mapping between local meta-data to optimal LoRA (Hu et al., 2022) ranks. Fig. 1(a) illustrates the personalization process of FedP2EFT on a single client during inference, where per-layer LoRA ranks are generated depending on the initial base model, the client’s dataset, and their resource budget. These personalized LoRA modules are then used to apply PEFT on the base model and yield the personalized model.
As FedP2EFT focuses on improving the PEFT process per-dataset/task/client, it is directly pluggable to any starting base model, which may or may not be federatedly trained. This includes off-the-shelf pretrained models, FL approaches that learn a single global model (McMahan et al., 2017; Oh et al., 2022; Karimireddy et al., 2020; Acar et al., 2021), and even personalized FL approaches that deploy personalized layers, e.g., monolingual tokenizer and embeddings (Iacob et al., 2025), personalized LoRA adapters (Qi et al., 2024; Wu et al., 2024; Yang et al., 2024b), and language embeddings (Silva et al., 2023). Through our experiments (Section 4), we show that our method 1) largely outperforms both existing non-FL LoRA rank selection and FL-based learning-to-personalize techniques, and 2) complements well with a range of existing FL approaches.
@@ 2 Related Work
Multilingual LLMs (MLLMs). Existing efforts in multilingual LLMs often underperform on low-resource languages due to 1) data scarcity (Xu et al., 2024b), 2) the model’s limited capacity to learn the intricacies of multiple languages (Conneau et al., 2020), and 3) negative transfer learning among languages (Wang et al., 2020b). Common ways to counteract these challenges include the use of separate vocabulary and embeddings (Artetxe et al., 2020), hand-crafted adapters (Pfeiffer et al., 2020), automatic data annotation (Dubey et al., 2024), clustering and merging languages with similar representations (Chung et al., 2020), among other contributions (Wang et al., 2020a; Conneau et al., 2019). Our work is orthogonal to these approaches and builds upon recent FL-based MLLMs (Zhao et al., 2024; Iacob et al., 2025; Ye et al., 2024), which utilize FL to tap into previously inaccessible low-resource data sources.
@@
Personalized Federated Learning. To obtain personalized client-specific models, various approaches have been proposed, including the use of personalized layers (Arivazhagan et al., 2019), meta-learning (Chen et al., 2018), model mixtures (Marfoq et al., 2021), hypernetworks (Shamsian et al., 2021), transfer learning between global and local models (Shen et al., 2020), among other contributions (Deng et al., 2020). Some of these techniques have also been adopted for LLMs, e.g., personalized LoRAs (Yang et al., 2024b), hypernetworks for client embeddings (Silva et al., 2023), and mixtures of LoRA experts (Zhang et al., 2024b). Our work complements these approaches as personalized models can benefit from further fine-tuning as shown in Section 4.2.1.
@@
Federated Hyperparameter Optimization (HPO). Most federated approaches to HPO do not utilize the client dataset for personalized hyperparameters. Instead, they employ a single set of hyperparameters across all clients based on the local validation loss evaluated before FL (Zhou et al., 2023; Holly et al., 2022) or sample from federatedly learnt hyperparameter categorical distributions for each client (Khodak et al., 2021). An exception to this is FedL2P (Lee et al., 2023) which utilizes a PSG for personalized per-layer learning rates and batch normalization hyperparameters. We compare with FedL2P in our experiments.
@@
PEFT Structure Learning. Contrary to the conventional approach of distributing uniform adapter modules across all layers, a recent line of work allows different LoRA ranks to be used across a model’s weight matrices. Using fine-grained per-layer rank selection, existing methods include SVD-based LoRA reformulation followed by importance-based rank assignment (Zhang et al., 2023), trainable rank-gating units (Ding et al., 2023), selectively employing parallel weight modules (Song et al., 2024), meta-learning-based (Zhang et al., 2024a) and black-box optimization techniques (Tribes et al., 2024), specialized training recipes for multi-rank LoRA modules that allow flexible extraction of a range of ranks (Valipour et al., 2023), and coarse-grained dropping of LoRA-enhanced layers (Yao et al., 2024). While these methods can be effective in centralized setups, they typically require an excessive number of optimization steps, which is prone to overfitting in FL settings, where clients have limited amount of data.
@@ 3.1 Preliminaries & Motivation
In personalized FL, the goal is to minimize each client’s local objective $\mathbb{E}_{(x,y)\sim P^{i}}\mathcal{L}^{i}(\Phi^{i};x,y)$ where $P^{i}$ represents the data distribution of the $i$-th client, $x$ and $y$ are the input data and labels, respectively, and $\mathcal{L}^{i}(\Phi^{i};x,y)$ is the loss function for client $i$ given model parameters $\Phi^{i}$. This is typically achieved via fine-tuning (Matsuda et al., 2022; Chen et al., 2022) a base model, with parameters $\Phi^{i}_{BM}$ and a set of hyperparameters, e.g. learning rate. Note that $\Phi^{i}_{BM}$ may differ across clients if it is already personalized, e.g. if $\Phi^{i}_{BM}$ is obtained using a personalized FL algorithm.
Fine-tuning LLMs, however, is unprecedentedly compute and memory intensive, and prone to overfitting. As such, the majority of existing federated LLM works (Zhao et al., 2024; Sun et al., 2024) rely on PEFT methods, with LoRA (Hu et al., 2022) being a prevalent choice due to its efficiency and performance. Specifically, for a frozen weight matrix $W\in\mathbb{R}^{d\times e}$, LoRA introduces low-rank matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times e}$ where $r\ll\min(d,e)$. The adapted weights are then expressed as: $W^{\prime}=W+\frac{\alpha_{\text{lora}}}{r}BA$ where $\alpha_{\text{lora}}$ is a hyperparameter and only $B$ and $A$ are trained during fine-tuning. Although effective, these FL works rely on a fixed hand-crafted PS, e.g., a manually defined LoRA rank on hand-picked layers, for all clients, leading to suboptimal personalized models.
@@ 3.2 Personalized PEFT
We, instead, propose using a different PS for each client. Common hyperparameter choices from previous federated HPO approaches (Section 2) include learning rates and batch normalization (BN) hyperparameters. While these hyperparameters have been shown to be effective for handling data heterogeneity in popular vision and speech benchmarks (Li et al., 2021, 2016; Arivazhagan et al., 2019), they are less consequential or not applicable when fine-tuning LLMs. This stems from the fact that LLMs are often fine-tuned using adaptive optimizers, e.g. Adam, which are more robust to the learning rate (Zhao et al., 2025), and BN layers are not typically used. A more critical hyperparameter choice shown to be effective, especially for cross-lingual transfer learning (Pfeiffer et al., 2020), is the PEFT adapter structure; specifically which layers to introduce LoRAs in and what ranks to utilize (Zhang et al., 2023, 2024a).
Adapting BayesTune for LoRA Rank Selection.  Building upon BayesTune (Kim & Hospedales, 2023), a Bayesian sparse model selection approach, we formulate PEFT personalization as a sparse LoRA rank selection problem and propose BayesTune-LoRA. Concretely, we introduce rank-wise latent variables $\lambda\in\mathbb{R}^{r},\;\lambda_{i}>0,\;\forall i=1,2,\cdots,r$ for each LoRA matrix: $B\lambda A$. Let $\bm{\lambda}=\{\lambda_{l,\cdot}\}_{l=1}^{L}$ be the set of all $\lambda$ where $\lambda_{l,\cdot}$ represents the rank-wise scales for layer $l$ in a model with $L$ LoRA modules (similarly for $\bm{A}$ and $\bm{B}$). Using BayesTune, the values for $\theta=(\bm{\lambda}$,$\bm{A}$,$\bm{B})$ are optimized as:
@@
where $D$$=$$\{(x_{i},y_{i})\}_{i=1}^{N}$ is the train dataset, $N$ the size of $D$, $\mathcal{L}_{\text{CE}}(\theta;D)$ the cross-entropy loss, $\alpha_{p}$ and $\alpha_{s}$ hyperparameters, $\mathcal{L}_{s}$ the logarithm of the Laplace distribution (prior imposed on $p(B|\lambda)$ [^1] Unlike BayesTune, where every parameter is associated with its own prior scale, we use an “independent” Laplace prior where each $\lambda_{l,i}$ applies to all entries of $B_{l,i}$), $f(\|B_{l,i}\|_{1};\mu,b)=\frac{1}{2b}\exp\left(-\frac{|\|B_{l,i}\|_{1}-\mu|}{b}\right)$ with $\mu=0$ ($B$ is initialized to 0 in LoRA) and $b=\lambda_{l,i}$:
@@
and $\mathcal{L}_{p}$ is the logarithm of the Gamma distribution (hyper-prior imposed on $\lambda$), $\mathcal{G}(\lambda_{l,i};\alpha_{g},\beta_{g})=\frac{\beta_{g}^{\alpha_{g}}}{\Gamma(\alpha_{g})}\lambda_{l,i}^{\alpha_{g}-1}e^{-\beta_{g}\lambda_{l,i}}$ where $\alpha_{g}=0.01,\beta_{g}=100$ following the hyperparameters set by the original authors:
@@
Roughly speaking, $\mathcal{L}_{p}$ encourages small $\lambda$ while $\mathcal{L}_{s}$ encourages larger $\lambda$ for larger LoRA $B$ (per column) updates. Hence, minimizing the losses in Eq. (1) encourages larger $\lambda$ in more significant ranks.
Personalizing PEFT with BayesTune-LoRA. For each client, we attach BayesTune-LoRA modules, $\theta$, to all linear layers of its base model with rank $r_{\text{init}}=\alpha_{r\_mul}\cdot r_{\text{max target}}$ where $r_{\text{max target}}$ is the maximum inference resource budget and $r_{\text{init}}$ is the initial rank before pruning. $\theta$ is then optimized using Adam (Kingma & Ba, 2015) as per Eq. (1). [^2] BayesTune proposed using SGLD (Welling & Teh, 2011), adding Gaussian noise to the gradient updates and sampling from the posterior distribution. Due to the challenges of estimating the full posterior distribution in FL settings, particularly with limited client data, we opt to find a point estimate.
After training, we freeze the resulting $\bm{\lambda}$ and use it for personalization. Specifically, given a resource budget (total rank budget) of $r\cdot L$, we prune $\bm{\lambda}$ by taking the top-$(r\cdot L)$ largest ranks, along with the corresponding rows of $\bm{A}$ and columns of $\bm{B}$ [^3] The LoRA module is discarded for layers where $\|\lambda_{l}\|_{1}=0$ We then reinitialize the pruned $\bm{A}$ and $\bm{B}$ and perform standard fine-tuning on $\mathcal{L}_{\text{CE}}$ with the frozen pruned $\bm{\lambda}$ to obtain the personalized model. Note that we only have to train $\bm{\lambda}$ once for all ranks $\leq r_{\text{max target}}$.

### v2

# FedP2EFT: Federated Learning to Personalize PEFT for Multilingual LLMs

@@ Abstract
Federated learning (FL) has enabled training of multilingual large language models (LLMs) on diverse and decentralized multilingual data, especially on low-resource languages. To improve client-specific performance, personalization via the use of parameter-efficient fine-tuning (PEFT) modules such as LoRA is common. This involves a personalization strategy (PS), such as the design of the PEFT adapter structures (e.g., in which layers to add LoRAs and what ranks) and choice of hyperparameters (e.g., learning rates) for fine-tuning. Instead of manual PS configuration, we propose FedP2EFT, a federated learning-to-personalize method for multilingual LLMs in cross-device FL settings. Unlike most existing PEFT structure selection methods, which are prone to overfitting low-data regimes, FedP2EFT collaboratively learns the optimal personalized PEFT structure for each client via Bayesian sparse rank selection. Evaluations on both simulated and real-world multilingual FL benchmarks demonstrate that FedP2EFT largely outperforms existing personalized fine-tuning methods, while complementing other existing FL methods. Code is available at https://github.com/SamsungLabs/fedp2eft.
@@ 1 Introduction
Federated learning (FL) makes it possible to train multilingual large language models (LLMs) across different geographical regions, protecting linguistic diversity for low-resource languages (Zhao et al. 2024) while being compliant with privacy regulations (Lim et al. 2020), e.g., General Data Protection Regulation (GDPR). Despite the impressive capabilities demonstrated by these models across various languages, their performance varies significantly depending on the language (Rust et al. 2021) and the data volume (Adelani et al. 2021) per client.
Moreover, the majority of existing FL-based multilingual LLM approaches have thus far focused on training a single global model (Iacob et al. 2025; Weller et al. 2022; Ye et al. 2024), limiting their performance on specific languages. Concretely, scaling a single model to different languages is challenged by issues such as the Curse of Multilinguality (Conneau et al. 2020), where adding more languages often leads to diminishing returns, and Negative Interference (Wang, Lipton, and Tsvetkov 2020), where diverse languages compete for limited model capacity. From a personalized FL (pFL) perspective, learning a global model often increases the initial performance at the expense of personalized performance, e.g., fine-tuning from the global model (Jiang et al. 2019).
Naturally, pFL approaches can help bridge the gap and improve language personalization. However, existing techniques are either too costly to be applied to LLMs, e.g., the use of meta-learning (Chen et al. 2018) and hypernetworks (Shamsian et al. 2021), or rely on suboptimal hand-crafted personalization strategies, e.g., manual choice of personalized and language-specific layers (Iacob et al. 2025; Zhao et al. 2024; Wu et al. 2024; Qi et al. 2024), parameter-efficient fine-tuning (PEFT) adapter structures (Yang et al. 2024a, b), or clustering based on class labels and/or language commonality (Mansour et al. 2020; Sattler, Müller, and Samek 2020; Ye et al. 2024).
Intuitively, optimizing personalization in pFL often necessitates dataset- and task-specific methods. The optimal level of personalization varies significantly depending on the characteristics of the data and the specific FL scenario (Chen et al. 2022; Ye et al. 2024). For instance, an English-pretrained LLM may require stronger personalization, e.g., higher learning rates, when fine-tuning on German than on English. The optimal personalization strategy (PS) is thus contingent upon the specific task, the client, and the given base model (Lee et al. 2023).
In this paper, we address the issues above by proposing a novel federated hyperoptimization strategy that learns personalized PEFT configurations for each client. To this end, we propose FedP2EFT, a method that enables clients to collaboratively learn language personalization strategies using FL. Specifically, we federatedly train a PS generator (PSG), as depicted in Fig. 1(a), which allows all participating clients to collaboratively learn the optimal mapping between local meta-data to optimal LoRA (Hu et al. 2022) ranks. Fig. 1(b) illustrates the personalization process of FedP2EFT on a single client during inference, where per-layer LoRA ranks are generated depending on the initial base model, the client’s dataset, and their resource budget. These personalized LoRA modules are then used to apply PEFT on the base model and yield the personalized model.
As FedP2EFT focuses on improving the PEFT process per-dataset/task/client, it is directly pluggable to any starting base model, which may or may not be federatedly trained. This includes off-the-shelf pretrained models, FL approaches that learn a single global model (McMahan et al. 2017; Oh, Kim, and Yun 2022; Karimireddy et al. 2020; Acar et al. 2021), and even pFL approaches that deploy personalized layers, e.g., monolingual tokenizer and embeddings (Iacob et al. 2025), personalized LoRA adapters (Qi et al. 2024; Wu et al. 2024; Yang et al. 2024b), and language embeddings (Silva, Tambwekar, and Gombolay 2023). Through our experiments (Section 4), we show that our method 1) largely outperforms both existing non-FL LoRA rank selection and FL-based learning-to-personalize techniques, and 2) complements well with a range of existing FL approaches.
@@
Figure 1: (a) We train our personalization strategy generator (PSG) using standard FL approaches. (b) FedP2EFT’s inference stage on a single client. 1 Given the base model and the client’s train dataset, features are extracted and passed into our PSG to generate a PS, $\bm{\lambda}$, suited to the client’s resource budget. 2 $\bm{\lambda}$ is then used to initialize all LoRA modules before 3 the base model is personalized. The resulting personalized model is then used to evaluate on the client’s test samples.
@@ 2 Related Work
Multilingual LLMs (MLLMs). Existing efforts in multilingual LLMs often underperform on low-resource languages due to 1) data scarcity (Xu et al. 2024b), 2) the model’s limited capacity to learn the intricacies of multiple languages (Conneau et al. 2020), and 3) negative transfer learning among languages (Wang, Lipton, and Tsvetkov 2020). Common ways to counteract these challenges include the use of separate vocabulary and embeddings (Artetxe, Ruder, and Yogatama 2020), hand-crafted adapters (Pfeiffer et al. 2020), automatic data annotation (Dubey et al. 2024), clustering and merging languages with similar representations (Chung et al. 2020), among other contributions (Wang et al. 2020; Conneau et al. 2019). Our work is orthogonal to these approaches and builds upon recent FL-based MLLMs (Zhao et al. 2024; Iacob et al. 2025; Ye et al. 2024), which utilize FL to tap into previously inaccessible low-resource data sources.
@@
Personalized Federated Learning (pFL). To obtain personalized client-specific models, various approaches have been proposed, including the use of personalized layers (Arivazhagan et al. 2019), meta-learning (Chen et al. 2018), model mixtures (Marfoq et al. 2021), hypernetworks (Shamsian et al. 2021), transfer learning between global and local models (Shen et al. 2020), among other contributions (Deng, Kamani, and Mahdavi 2020). Some of these techniques have also been adopted for LLMs, e.g., personalized LoRAs (Yang et al. 2024b), hypernetworks for client embeddings (Silva, Tambwekar, and Gombolay 2023), and mixtures of LoRA experts (Zhang et al. 2024b). Our work complements these approaches as personalized models can benefit from further fine-tuning as shown in Section 4.4.
@@
Federated Hyperparameter Optimization (HPO). Most federated approaches to HPO do not utilize the client dataset for personalized hyperparameters. Instead, they employ a single set of hyperparameters across all clients based on the local validation loss evaluated before FL (Zhou et al. 2023; Holly et al. 2022) or sample from federatedly learnt hyperparameter categorical distributions for each client (Khodak et al. 2021). An exception to this is FedL2P (Lee et al. 2023) which utilizes a PSG for personalized per-layer learning rates and batch normalization hyperparameters. While FedL2P has been shown to work well in standard small-scale image and speech benchmarks, their applicability to LLMs is unclear: i) LLMs don’t use batch normalization (BN), ii) LLMs often use adaptive optimizers, making learning rate a less sensitive hyparam for downstream performance. Learning the learning rates also require FedL2P to adopt expensive 2nd-order optimization methods as the learning rate is not a direct gradient of the loss, a limitation that is further aggravated with LLMs. We compare with FedL2P in our experiments.
@@
PEFT Structure Learning. Contrary to the conventional approach of distributing uniform adapter modules across all layers, a recent line of work allows different LoRA ranks to be used across a model’s weight matrices. Using fine-grained per-layer rank selection, existing methods include SVD-based LoRA reformulation followed by importance-based rank assignment (Zhang et al. 2023), trainable rank-gating units (Ding et al. 2023), selectively employing parallel weight modules (Song et al. 2024), meta-learning-based (Zhang et al. 2024a) and black-box optimization techniques (Tribes et al. 2024), specialized training recipes for multi-rank LoRA modules that allow flexible extraction of a range of ranks (Valipour et al. 2023), and coarse-grained dropping of LoRA-enhanced layers (Yao et al. 2024). While these methods can be effective in centralized setups, they typically require an excessive number of optimization steps, which is prone to overfitting in FL settings, where clients have limited amount of data.
@@ 3.1 Preliminaries & Motivation
In pFL, the goal is to minimize each client’s local objective $\mathbb{E}_{(x,y)\sim P^{i}}\mathcal{L}^{i}(\Phi^{i};x,y)$ where $P^{i}$ represents the data distribution of the $i$-th client, $x$ and $y$ are the input data and labels, respectively, and $\mathcal{L}^{i}(\Phi^{i};x,y)$ is the loss function for client $i$ given model parameters $\Phi^{i}$. This is typically achieved via fine-tuning (Matsuda et al. 2022; Chen et al. 2022) a base model, with parameters $\Phi^{i}_{BM}$ and a set of hyperparameters, e.g. learning rate. Note that $\Phi^{i}_{BM}$ may differ across clients if it is already personalized, e.g. if $\Phi^{i}_{BM}$ is obtained using a pFL algorithm.
Fine-tuning LLMs, however, is unprecedentedly compute and memory intensive, and prone to overfitting. As such, the majority of existing federated LLM works (Zhao et al. 2024; Sun et al. 2024) rely on PEFT methods, with LoRA (Hu et al. 2022) being a prevalent choice due to its efficiency and performance. Specifically, for a frozen weight matrix $W\in\mathbb{R}^{d\times e}$, LoRA introduces low-rank matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times e}$ where $r\ll\min(d,e)$. The adapted weights are then expressed as: $W^{\prime}=W+\frac{\alpha_{\text{lora}}}{r}BA$ where $\alpha_{\text{lora}}$ is a hyperparameter and only $B$ and $A$ are trained during fine-tuning. Although effective, these FL works rely on a fixed hand-crafted PS, e.g., a manually defined LoRA rank on hand-picked layers, for all clients, leading to suboptimal personalized models.
@@ 3.2 Personalized PEFT
We, instead, propose using a different PS for each client. Common hyperparameter choices from previous federated HPO approaches (Section 2) include learning rates and BN hyperparameters. While these hyperparameters have been shown to be effective for handling data heterogeneity in popular vision and speech benchmarks (Li et al. 2021, 2016; Arivazhagan et al. 2019), they are less consequential or not applicable when fine-tuning LLMs. This stems from the fact that LLMs are often fine-tuned using adaptive optimizers, e.g. Adam, which are more robust to the learning rate (Zhao et al. 2025), and BN layers are not typically used. A more critical hyperparameter choice shown to be effective, especially for cross-lingual transfer learning (Pfeiffer et al. 2020), is the PEFT adapter structure; specifically which layers to introduce LoRAs in and what ranks to utilize (Zhang et al. 2023, 2024a).
Adapting BayesTune for LoRA Rank Selection. We build upon BayesTune (Kim and Hospedales 2023), a Bayesian sparse model selection approach. Directly using BayesTune in our approach, however, will not only incur an intractable cost but also be difficult to optimize due to the high dimensionality of predicting per-parameter scaling coefficients (Appendix C). Hence, we adapt BayesTune by formulating PEFT personalization as a sparse LoRA rank selection problem and propose BayesTune-LoRA (BT-LoRA). Concretely, we introduce rank-wise latent variables $\lambda\in\mathbb{R}^{r},\;\lambda_{i}>0,\;\forall i=1,2,\cdots,r$ for each LoRA matrix: $B\lambda A$. Let $\bm{\lambda}=\{\lambda_{l,\cdot}\}_{l=1}^{L}$ be the set of all $\lambda$ where $\lambda_{l,\cdot}$ represents the rank-wise scales for layer $l$ in a model with $L$ LoRA modules (similarly for $\bm{A}$ and $\bm{B}$). Using BayesTune, the values for $\theta=(\bm{\lambda}$$\bm{A}$$\bm{B})$ are optimized as:
@@
[…]
@@
[…]
@@
Intuition of $\mathcal{L}_{p}$ and $\mathcal{L}_{s}$. $\mathcal{L}_{p}$ encourages small $\lambda$ while $\mathcal{L}_{s}$ encourages larger $\lambda$ for larger LoRA $B$ (per column) updates. Hence, minimizing the losses in Eq. (1) encourages larger $\lambda$ in more significant ranks.
Personalizing PEFT with BT-LoRA. For each client, we attach BT-LoRA modules, $\theta$, to all linear layers of its base model with rank $r_{\text{init}}=\alpha_{r\_mul}\cdot r_{\text{max target}}$ where $r_{\text{max target}}$ is the maximum inference resource budget and $r_{\text{init}}$ is the initial rank before pruning. $\theta$ is then optimized using Adam (Kingma and Ba 2015) as per Eq. (1). [^3] BayesTune proposed using SGLD (Welling and Teh 2011), adding Gaussian noise to the gradient updates and sampling from the posterior distribution. Due to the challenges of estimating the full posterior distribution in FL settings, particularly with limited client data, we opt to find a point estimate.
After training, we freeze the resulting $\bm{\lambda}$ and use it for personalization. Specifically, given a resource budget (total rank budget) of $r\cdot L$, we prune $\bm{\lambda}$ by taking the top-$(r\cdot L)$ largest ranks, along with the corresponding rows of $\bm{A}$ and columns of $\bm{B}$. [^4] The LoRA module is discarded for layers where $\|\lambda_{l}\|_{1}=0$ We then reinitialize the pruned $\bm{A}$ and $\bm{B}$ and perform standard fine-tuning on $\mathcal{L}_{\text{CE}}$ with the frozen pruned $\bm{\lambda}$ to obtain the personalized model. Note that we only have to train $\bm{\lambda}$ once for all ranks $\leq r_{\text{max target}}$.